One packet of biscuits requires 2½ cups of flour and 1 whole 2/3 cups of sugar. Estimated total quantity of both ingredients used in 10 such packets of biscuits will be (a) less than 30 cups (b) between 30 cups and 40 cups (c) between 40 cups and 50 cups (d) above 50 cups

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
From the question it is given that,
One packet of biscuits requires 21⁄2 cups of flour = 5/2

One packet of biscuits requires cups of sugar. = 5/3 Total ingredients for one packet of biscuits = (5/2) + (5/3)

= (15 + 10)/6

= 25/6
Then, total quantity of both ingredients used in 10 such packets of biscuits = 10 × (25/6)

= 5 × (25/3)
= 125/3

Thanks